# Bridging Gaps in Natural Language Processing for Yorùbá: A Systematic Review of a Decade of Progress and Prospects

### v1

@@ 4.1 General Statistics
The primary studies included in the research, totalling $105$, were synthesized for relevant data. They consist only of journal articles and conference papers based on the selection criteria specified in the systematic review protocol. Figure 6 reveals the frequency of the sample—included primary studies—over the years. It attests that the Yorùbá NLP research has grown over the years, with an apparent upward trend after the break in year $2017$. In addition, $2023$ and $2024$ have the highest number of primary studies, even though the latter year is yet to end at the time of this review. Subsequently, the publications are assessed based on the type. Figure 7 shows the distribution of article type, in which there are more conference papers $(62.9\%)$ than journal articles $(37.1\%)$. This observation is plausible as most researchers in this domain are primarily involved in conferences.
Figure 6: Distribution of publication per year
@@
Furthermore, accessing the authors contributing to this field shows an informative trend, as depicted in Figure 8. Figure 8(b) shows the top 20 authors in this field based on the primary studies. The top author, David Adelani, has about 18 included publications in total. Moreover, there have been observable collaborations among the top $7$ authors, including Dossou, BFP; and Osei, S. Similarly, significant collaborations exist among Lin, J; Oladipo, A; and Ogundepo, OJ, among other notable collaborations. In addition, even though the top $20$ authors in the field were represented, the minimum number of publications on the chart in Figure 8(b) being $5$, left out authors with publications between $2$ and $4$. Hence, Figure 8(a) ensures a representative visualization by revealing famous authors in the research domain.
(a) Common authors
(b) Top 20 authors
@@
Finally, Figure 9 shows the first authors’ geographical locations at the publication’s time. It reveals an interesting insight, making it known that most publishers were based in Nigeria at the time of publishing, while a considerable number were also in the United States, Canada, South Africa, Germany, and the Benin Republic. This tells that research in this domain extends beyond the dominant speaking country of the Language—Nigeria. It also shows promising outcomes due to visible collaborations taking place among researchers across the world.
Figure 9: Publications by geographical location
@@ 4.2 Research Question Analysis
4.2.1 RQ1: What NLP tasks have been addressed for the Yorùbá language?
@@
Efforts have been made to democratize the numbering system of Yorùbá language through NLP with an earlier study [49], which tried to develop a computational system for converting cardinal numbers into their Yorùbá number names in a number-to-text conversion task. Similarly, [50] develops a machine translation system for translating English number text into Yorùbá text, considering the context. While the former focused on the text normalization task, a core foundational NLP task required in this research domain, the latter focused on machine translation. Moreover, diacritic restoration, an important component of text normalization tasks, has been given considerable research efforts, as seen from [31, 51, 52, 53] studies. Ultimately, three studies [54, 55, 56] have made efforts toward improving spell-checking and correction in the Yorùbá texts. These tasks are crucial for accurately representing texts in the language, thereby improving the availability of quality data.
@@
Many studies also explored the crucial task involving automatic translation of text data between Yorùbá and other languages. About $18$ studies explored the machine translation (MT) tasks, cutting across rule-based machine translation (RBMT) [30, 64, 65, 66, 67, 50, 9], statistical machine translation (SMT) [8, 68], neural machine translation (NMT) [28, 10, 69, 70, 71, 72], and a few hybrid MT methods involving SMT and RBMT [29]; RBMT and example-based machine translation (EBMT) [73]; SMT and NMT [7]. Moreover, a similar task on sentence alignment, which is crucial to NMT, was explored in [74], considering a bilingual approach involving English-Yorùbá pair.
@@
Even though most studies tend to involve text processing, quite a substantial number of studies also investigated speech processing in Yorùbá NLP. This ranges from text-to-speech synthesis [87, 88, 89, 90], tone recognition for continuous speech [91], speech recognition [92, 93, 94], text-to-speech analysis [95], speech-based gender recognition [96] and multi-tasks involving both text-to-speech and speech-to-text [97], text-to-speech and language speech resource development [98, 99, 36, 100, 101], speech synthesis and language pitch modelling [102], and two studies exclusively focusing on speech corpus development [103, 104]. Whereas, [35] differs as it is based on corpus development for visually grounded speech. Furthermore, to optimize modelling in speech recognition, acoustic unit discovery tasks [105, 106, 107] are essential in learning speech sounds embedding to retain the linguistically relevant acoustic information and discard the irrelevant ones.
Ultimately, increased usage of social networking sites and websites, among others, has undoubtedly exploded the extent of information available in recent years. Thus, developing effective methods of accessing and retrieving information is indispensable. Few studies also explored this, including tasks on dense information retrieval [108], monolingual retrieval tasks [109], and cross-lingual information retrieval tasks [18, 110, 111].
@@
Language modelling [38], [75], [77], [76] Multilingual pre-training WÚRÀ; Common Crawl Corpus; Custom data
@@
Machine translation [8], [29], [30], [28], [71], [9], [64], [70], [69], [65], [7], [73], [66], [67], [74], [50], [68], [72], [10], [34], [112], [37] Rule-based; statistical modelling; multilingual pre-training WebCrawl African; YORÙLECT; IkiniYorùbá; MENYO-20k; Custom data
@@
Text and topic classification [115], [83], [82], [77], [84], [85], [86] Machine learning; multilingual language pre-training MasakhaNEWS; Common Crawl Corpus; Custom data
@@
4.2.2 RQ2: What techniques have been employed for Yorùbá NLP?
The metamorphosis of NLP techniques involving Yorùbá language is observed to not necessarily lag behind the general trend of accomplishment in the NLP community. Most of the earlier research works are seen to be limited to implementing rule-based techniques [87, 30, 64, 55] before the advancement, leading to incorporating statistical modelling [103, 86, 8, 116], and sometimes, a combination of both techniques [81]. Furthermore, recent studies have incorporated supervised machine learning [11, 51, 82, 93] and unsupervised learning [117, 107, 57, 105, 59] techniques in this research domain, and a few times, the hybrid of statistics and machine learning techniques [33, 116]. Lately, deep learning architectures [80, 79, 52, 31, 28], as well as a hybrid of machine and deep learning methods [115] are being utilised for various tasks as well.
Also, due to the low-resource nature of Yorùbá language, NLP research works in its domain have benefitted from transfer learning by pre-training on a high-resource language and fine-tuning on specific tasks in this language. Transfer learning of this nature over the years has been mainly composed of leveraging multilingual pre-training [71, 118, 119, 120, 63, 77] and cross-lingual transfer [121, 18, 17, 111]. Figure 10 shows word clouds of both the employed techniques 10(b) and the NLP tasks 10(a) carried out in the primary studies involved in the review.
(a) Common NLP tasks
(b) Common NLP techniques
Figure 10: Word clouds of NLP tasks and techniques
Table LABEL:tab6 shows a cross-tabulation of the specific NLP tasks involving Yorùbá language together with the NLP techniques utilised in each task. It reveals that only studies involving sentiment analysis and TTS tasks utilised all the techniques as classified. The MT task closely followed the previous two tasks vis-a-vis the varieties of techniques involved in their studies; all the techniques classes were utilised in the MT studies, except for the machine learning technique. Moreover, other tasks have studies utilising most of the classes of techniques; however, acoustic unit discovery, ASR, spell checking & correction, speech-to-text, and tone identification & recognition tasks include studies using only two different techniques. In addition, only the speech-based gender recognition task has studies involving only a single technique—deep learning technique.
@@ 4.2.3 RQ3: What language resources are available for Yorùbá language?
Many low-resource languages experience stunted growth in their language development due to the flaw associated with their lack of substantial datasets required for various NLP tasks amidst recent data-hungry models. This factor has undoubtedly prompted significant efforts at developing various language resources for Yorùbá language to further improve the performance of various NLP tasks in the language. The developed resources have been seen to cut across different domains; while some studies solely focused on resource development, others utilised the datasets or corpora on specific tasks. The summary of the corpora, which contains Yorùbá either as a monolingual, bilingual, or as an important part of a multilingual dataset, is presented in Table LABEL:tab7. Furthermore, for detail purposes, specific datasets or corpora with assigned names are briefly outlined individually as follows, while others are listed as part of the general descriptions:
AFRIQA: AFRIQA [18] pioneers among cross-lingual question-answering (QA) corpus development for African languages. It is the first cross-lingual open retrieval question-answering (XOR QA) dataset focused on African languages, bridging the gap in resource availability and paving the way for more equitable and inclusive question-answering technologies. More importantly, Yorùbá language is a core component among the languages in the dataset, as it comprises about $12,239$ XOR QA examples for $10$ African languages across the southern, western, eastern and central African regions, viz: Bemba, Fon, Hausa, Igbo, Kinyarwanda, Swahili, Twi, Wolof, Yorùbá and Zulu. Specifically, for Yorùbá language, the training, dev, and testing examples are $360,261$, and $332$, respectively. Furthermore, experiments to evaluate the performance of state-of-the-art models on the dataset were carried out, and this mainly involved automatic translation, which entails translating questions or retrieved documents in preparation for processing, and multilingual retrieval, involving carrying out direct retrieval through the aid of multilingual embeddings.
@@
AfriSenti: AfriSenti [4] focuses on providing a high-quality, large-scale dataset to address the lack of resources for sentiment analysis in underrepresented African languages. This benchmark contains over $110,000$ posts involving $14$ African languages across four language families—Niger-Congo, Afro-Asiatic, English Creole and Indo-European—in which Yorùbá is a core part. The languages featured in the dataset are namely: Amharic, Algerian Arabic, Hausa, Igbo, Kinyarwanda, Moroccan Arabic, Mozambican Portuguese, Nigerian Pidgin, Oromo, Swahili, Tigrinya, Twi, Xitsonga, and Yorùbá. The sentiment categories captured in the corpus include positive, negative and neutral. Furthermore, baseline traditional machine learning and pre-trained multilingual language models were trained to facilitate empirical analysis. Moreover, evaluation metrics include accuracy, F1 score, and precision/recall. Overall, it established a benchmark for sentiment analysis in Yorùbá language, also enabling comparisons of models and techniques for African languages.
@@
AfriWOZ: AfriWOZ [122] also pioneers in resource development involving dialogue generation for low-resource African languages. It presents a high-quality dialogue generation dataset for $6$ African languages, emphasising the Yorùbá language as a subset. Other five languages include Hausa, Kinyarwanda, Nigerian Pidgin English, Swahili, and Wolof; all languages spread across three language families—Afro-Asiatic, Niger-Congo, and English Creole. The dialogue dataset for the included languages was developed from the MultiWoz dataset [123], and experiments were performed to avail empirical analysis for related tasks using transfer learning approaches.
@@
Antenatal Orientation Speech Dataset: This dataset contains multilingual speech datasets in English, Nigerian Pidgin English, and Yorùbá languages, basically for antenatal orientation in Nigeria, being the main geographical source of the languages. The dataset [100] is composed of a word count of $2,639$ for English, $3,202$ for Yorùbá, and $2,521$ for Pidgin. Moreover, the corresponding speech datasets had $59,880$ seconds for English, $70,380$ seconds for Yorùbá, and $69,840$ seconds for Pidgin, both showing a dominant percentage for Yorùbá language. The size of the speech datasets was $15.6$ KB for English, $17.9$ KB for Yorùbá, and $18.3$ KB for Pidgin. This dataset will be valuable for antenatal orientation in Nigeria and contribute to the currently low-resource status of Yorùbá language by improving its availability for NLP tasks.
@@
BibleTTS: BibleTTS [98] is an open speech multilingual dataset containing ten languages of sub-Saharan Africa, viz: Akuapem Twi, Asante Twi, Chichewa, Ewe, Hausa, Kikuyu, Lingala, Luganda, Luo, and Yorùbá. The corpus emanates from Bible recordings provided by the Open-Bible project of Biblica and contains high-quality $48kHz$ studio recordings for single speakers, with up to $86$ hours of data per language. Furthermore, TTS models were developed by leveraging the dataset, which ascertains its usability and quality for speech synthesis. The dataset is openly available for commercial and non-commercial use as it’s licensed CC-BY-SA License, promoting accessibility and further research in text-to-speech for Yorùbá, among other languages captured.
@@
IkiniYorùbá: IkiniYorùbá[37] is a dataset focusing on the cultural aspect of Yorùbá, intending to investigate how well MT models can translate greetings in the language into English. This dataset comprises $960$ parallel sentences of Yorùbá greetings and their English translations. Moreover, it also incorporates the movie transcript subset of the MENYO-20k dataset[72], which contains conversational Yorùbá-English data. The performance of MT models was evaluated on the datasets to enable empirical analysis for related studies in future. The introduction of the IkiniYorùbá dataset is a valuable contribution to the research community working on Yorùbá NLP.
ÌròyìnSpeech: ÌròyìnSpeech [101] presents high-quality speech corpus exclusively in Yorùbá language, which allows for multipurpose usage involving Automatic Speech Recognition(ASR) and Text-to-Speech (TTS) tasks. The data collection involves curating 23,000 sentences from news and creative writing domains under an open license (CC-BY-4.0). Also, it included 5,000 sentences on the Mozilla Common Voice platform to crowdsource recordings and validations. The data contributions include 42 hours of in-house recorded speech data by 80 volunteers and an additional 6 hours of validated recordings on Mozilla Common Voice. Ultimately, a high-fidelity single-speaker Yorùbá TTS system was evaluated on the curated speech data, including a baseline word error rate (WER) of $23.8$ for ASR purposes.
@@
MasakhaPOS: MasakhaPOS [17] presents a huge POS dataset available for $20$ diverse African languages, of which Yorùbá language was greatly emphasised. Other languages featured in the dataset include Bambara, Fon, Hausa, Igbo, Luo, Luganda, Akan/Twi, Wolof, isiZulu, isiXhosa, Setswana, Kiswahili, Nigerian Pidgin English, Mossi, Kinyarwanda, Ghomálá, Éwé, chiShona, and Chichewa. Moreover, Yorùbá data was jointly obtained from Voice of Nigeria and Asejere online newspapers, containing a $43,601$ token with an average sentence length of $24.4$. Moreover, baseline models using conditional random fields (CRF) and multilingual pre-trained language models such as mBERT and XLM-R were developed to support empirical analysis for related tasks.
MENYO-20k: MENYO-20k [72] presents a multi-domain parallel corpus of Yorùbá-English with clean orthography and standardized train-test splits, with focus on improving the evaluations of MT models on low-resource language. The dataset’s domain includes texts from news articles, radio and movie transcripts, TED talks, etc. In addition, special attention was given to the diacritization of Yorùbá texts in the corpus, as it plays an essential role in intelligibility and translation quality in MT NLP tasks. Neural MT models are benchmarked to enable future empirical analysis in similar studies. The dataset is available through a CC BY-NC 4.0 licence.
@@
NollySenti: NollySenti [5] is a parallel multilingual corpus for sentiment classification in five languages spoken in Nigeria, including Yorùbá, Nigerian Pidgin English, Hausa, Igbo and English. The dataset source reviews from Nollywood movies—movies primarily made in Nigeria—initially in English, with translations into the four Nigerian languages involved. Moreover, the initial collection for the dataset comprises $882$ negative reviews and $1,018$ positive reviews, with an average word length of $20.7$ and $35.0$, respectively, making up two sentiment classification categories. Benchmarks were also developed through traditional machine-learning techniques and pre-trained language modelling to enable empirical analysis for future research in the domain.
Prosodic Read Speech Corpus: This corpus contains a high-quality speech corpus [104] of $7,376$ phrases and sentences in Standard Yorùbá language. A TTS system was developed on the corpus and evaluated using the mean open score and semantically unpredictable sentence (SUS) test score to support empirical analysis for future related studies. The developed corpus can be used as valuable research material for future studies on Yorùbá speech processing, synthesis, and recognition
TATA: TATA [119] dataset is a large and high-quality multilingual dataset focused on African languages, created from Demographic and Health Surveys (DHS) Program bilingual reports. It consists of $8,700$ examples in nine languages, with Russian as a zero-shot test language. Yorùbá is a core component of these languages; other languages in the dataset include Arabic, English, French, Hausa, Igbo, Portuguese, and Swahili. A transformer-based multilingual pre-trained language model was evaluated on the dataset, supporting empirical analysis in future related studies.

### v2

@@ 4.1 General Statistics
The primary studies included in the research, totalling $105$, were synthesized for relevant data. They consist only of journal articles and conference papers based on the selection criteria specified in the systematic review protocol. Figure 7 reveals the frequency of the sample—included primary studies—over the years. It attests that the Yorùbá NLP research has grown over the years, with an apparent upward trend after the break in year $2017$. In addition, $2023$ and $2024$ have the highest number of primary studies, even though the latter year is yet to end at the time of this review. Subsequently, the publications are assessed based on the type. Figure 7 shows the distribution of article type, in which there are more conference papers $(62.9\%)$ than journal articles $(37.1\%)$. This observation is plausible as most researchers in this domain are primarily involved in conferences, probably as conference publications are typically shorter and faster to publish.
Figure 6: Distribution of publications per year
@@
Furthermore, Figure 8 shows the first authors’ geographical locations at the publication’s time. It reveals an interesting insight, making it known that most publishers were based in Nigeria at the time of publishing, while a considerable number were also in the United States, Canada, South Africa, Germany, and the Benin Republic. This tells that research in this domain extends beyond the dominant speaking country of the Language—Nigeria. It also shows promising outcomes due to visible collaborations taking place among researchers across the world.
Figure 8: Authors’ geographical location
Finally, accessing the authors’ contributions to this field shows an informative trend, as depicted in Figure 9. It shows the top 20 authors in this field based on the primary studies. The top author, David Adelani, has about $18$ publications in total. Moreover, there have been observable collaborations among the top $7$ authors, including Dossou, BFP; and Osei, S. Similarly, significant collaborations exist among Lin, J; Oladipo, A; and Ogundepo, OJ, among other notable collaborations. This diagram reveals famous authors in the research domain.
@@ 4.2 Research Question Analysis
4.2.1 RQ1: What NLP tasks have been addressed for Yorùbá language?
@@
Efforts have been made to democratize the numbering system of Yorùbá language through NLP, with earlier studies by making digital content and technologies more inclusive and usable for native speakers. [28] tried to develop a computational system for converting cardinal numbers into their Yorùbá number names in a number-to-text conversion task, enabling machine-readable and voice-accessible numerical expressions in Yor‘ub’a. Similarly, [57] develops a machine translation system that translates English number expressions into Yorùbá text, taking into account the linguistic and syntactic context in which the numbers appear. While the former focused on the text normalization task, a core foundational NLP task required in this research domain, the latter focused on machine translation. Moreover, diacritic restoration, an important component of text normalization tasks, has been given considerable research efforts, as seen from [101, 44, 35, 34] studies. Ultimately, three studies [97, 92, 37] have made efforts toward improving spell-checking and correction in the Yorùbá texts. These tasks are crucial for accurately representing texts in the language, thereby improving the availability of quality data.
@@
Many studies also explored the crucial task involving automatic translation of text data between Yorùbá and other languages. About $18$ studies explored the machine translation (MT) tasks, cutting across rule-based machine translation (RBMT) [16, 108, 9, 67, 46, 57, 59], statistical machine translation (SMT) [41, 47], neural machine translation (NMT) [6, 23, 61, 80, 5, 13], and a few hybrid MT methods involving SMT and RBMT [62]; RBMT and example-based machine translation (EBMT) [15]; SMT and NMT [2]. Moreover, a similar task on sentence alignment, which is crucial to NMT, was explored in [112], considering a bilingual approach involving the English-Yorùbá pair.
@@
Even though most studies tend to involve text processing, quite a substantial number of studies also investigated speech processing in Yorùbá. This ranges from text-to-speech synthesis [70, 51, 36, 1], tone recognition for continuous speech [49], speech recognition [69, 115, 104], text-to-speech analysis [72], speech-based gender recognition [110] and multi-tasks involving both text-to-speech and speech-to-text [29], text-to-speech and language speech resource development [83, 53, 65, 24, 90], speech synthesis and language pitch modelling [123], and two studies exclusively focusing on speech corpus development [116, 30]. Whereas, [96] differs as it is based on corpus development for visually grounded speech. Furthermore, to optimize modelling in speech recognition, acoustic unit discovery tasks [99, 25, 127] are essential in learning speech sounds embedding to retain the linguistically relevant acoustic information and discard the irrelevant ones.
Ultimately, increased usage of social networking sites and websites, among others, has undoubtedly increased the extent of information available in recent years. Thus, developing effective methods of accessing and retrieving information is indispensable. Few studies also explored this, including tasks on dense information retrieval [94], monolingual retrieval tasks [128], and cross-lingual information retrieval tasks [89, 100, 19].
@@
4.2.2 RQ2: What techniques have been employed for Yorùbá NLP?
The metamorphosis of NLP techniques involving Yorùbá language is observed to not necessarily lag behind the general trend of accomplishment in the NLP community. Most of the earlier research works are seen to be limited to implementing rule-based techniques [70, 16, 108, 92] before the advancement, leading to incorporating statistical modelling [116, 39, 41, 98], and sometimes, a combination of both techniques [14]. Furthermore, recent studies have incorporated supervised machine learning [68, 44, 42, 115] and unsupervised learning [20, 127, 8, 99, 7] techniques in this research domain, and a few times, the hybrid of statistics and machine learning techniques [43, 98]. Lately, deep learning architectures [106, 105, 35, 101, 6], as well as a hybrid of machine and deep learning methods [95], are being utilised for various tasks as well.
Also, due to the low-resource nature of Yorùbá language, NLP research works in its domain have benefited from transfer learning by pre-training on a high-resource language and fine-tuning on specific tasks in this language. Transfer learning of this nature over the years has been mainly composed of leveraging multilingual pre-training [5, 3, 63, 107, 55, 88] and cross-lingual transfer [26, 89, 54, 19]. Figure 10 shows word clouds of both the employed techniques 10(b) and the NLP tasks 10(a) carried out in the primary studies involved in the SLR.
Table LABEL:tab6 shows a cross-tabulation of the specific NLP tasks involving Yorùbá language together with the NLP techniques utilised in each task. It reveals that only studies involving sentiment analysis and TTS tasks utilised all the techniques as classified. The MT task closely followed the previous two tasks vis-à-vis the varieties of techniques involved in their studies; all the techniques classes were utilised in the MT studies, except for the machine learning technique. Moreover, other tasks have studies utilising most of the classes of techniques; however, acoustic unit discovery, ASR, spell checking & correction, speech-to-text, and tone identification & recognition tasks include studies using only two different techniques. In addition, only the speech-based gender recognition task has studies involving only a single technique—a deep learning technique.
(a) Common NLP tasks
(b) Common NLP techniques
Figure 10: Word clouds of NLP tasks and techniques
Many low-resource languages experience stunted growth in their language development due to the flaw associated with their lack of substantial datasets required for various NLP tasks amidst recent data-hungry models. This factor has undoubtedly prompted significant efforts at developing various language resources for Yorùbá language to further improve the performance of various NLP tasks in the language. The developed resources have been seen to cut across different domains; while some studies solely focused on resource development, others utilised the datasets or corpora on specific tasks. The summary of the corpora, which contains Yorùbá either as a monolingual, bilingual, or as an important part of a multilingual dataset, is presented in Table LABEL:tab7. Furthermore, for detailed purposes, specific datasets or corpora with assigned names are briefly outlined individually as follows, while others are listed as part of the general descriptions.
AFRIQA: AFRIQA [89] pioneers among cross-lingual question-answering (QA) corpus development for African languages. It is the first cross-lingual open retrieval question-answering (XOR QA) dataset focused on African languages, bridging the gap in resource availability and paving the way for more equitable and inclusive question-answering technologies. More importantly, Yorùbá language is a core component among the languages in the dataset, as it comprises about $12,239$ XOR QA examples for $10$ African languages across the southern, western, eastern and central African regions, viz: Bemba, Fon, Hausa, Igbo, Kinyarwanda, Swahili, Twi, Wolof, Yorùbá and Zulu. Specifically, for the Yorùbá language, the training, development, and testing examples are $360,261$, and $332$, respectively. Furthermore, experiments to evaluate the performance of state-of-the-art models on the dataset were carried out, and this mainly involved automatic translation, entailing translating questions or retrieved documents in preparation for processing, and multilingual retrieval, involving carrying out direct retrieval through multilingual embeddings.
@@ 4.2.3 RQ3: What language resources are available for Yorùbá language?
Language modelling [91], [56], [88], [93], [75] Multilingual pre-training WÚRÀ; Common Crawl Corpus; Custom data
@@
Machine translation [41], [62], [16], [6], [5], [59], [108], [80], [61], [9], [2], [15], [67], [46], [112], [57], [47], [13], [23], [22], [124], [27] Rule-based; statistical modelling; multilingual pre-training WebCrawl African; YORÙLECT; IkiniYorùbá; MENYO-20k; Custom data,
@@
Text and topic classification [95], [11], [42], [88], [66] Machine learning; multilingual language pre-training MasakhaNEWS; Common Crawl Corpus; Custom data
@@
AfriSenti: AfriSenti [85] focuses on providing a high-quality, large-scale dataset to address the lack of resources for sentiment analysis in underrepresented African languages. This benchmark contains over $110,000$ posts involving $14$ African languages across four language families—Niger-Congo, Afro-Asiatic, English Creole and Indo-European—in which Yorùbá is a core part. The languages featured in the dataset are, namely: Amharic, Algerian Arabic, Hausa, Igbo, Kinyarwanda, Moroccan Arabic, Mozambican Portuguese, Nigerian Pidgin, Oromo, Swahili, Tigrinya, Twi, Xitsonga, and Yorùbá. The sentiment categories captured in the corpus include positive, negative, and neutral. Furthermore, baseline traditional machine learning and pre-trained multilingual language models were trained to facilitate empirical analysis. Moreover, evaluation metrics include accuracy, F1 score, and precision/recall. Overall, it established a benchmark for sentiment analysis in Yorùbá language, also enabling comparisons of models and techniques for African languages.
@@
AfriWOZ: AfriWOZ [18] also pioneers in resource development involving dialogue generation for low-resource African languages. It presents a high-quality dialogue generation dataset for $6$ African languages, emphasising the Yorùbá language as a subset. The other five languages include Hausa, Kinyarwanda, Nigerian Pidgin English, Swahili, and Wolof; all languages belong to three language families: Afro-Asiatic, Niger-Congo, and English Creole. The dialogue dataset for the included languages was developed from the MultiWoz dataset [52], and experiments were performed to provide empirical analysis for related tasks using transfer learning approaches.
@@
Antenatal Orientation Speech Dataset: This dataset comprises multilingual speech samples in English, Nigerian Pidgin English, and Yorùbá languages, primarily for antenatal health education in Nigeria, which is the primary geographical source of these languages. The dataset [24] is composed of a word count of $2,639$ for English, $3,202$ for Yorùbá, and $2,521$ for Pidgin. Moreover, the corresponding speech datasets had $59,880$ seconds for English, $70,380$ seconds for Yorùbá, and $69,840$ seconds for Pidgin, both showing a dominant percentage for the Yorùbá language. The size of the speech datasets was $15.6$ KB for English, $17.9$ KB for Yorùbá, and $18.3$ KB for Pidgin. This dataset will be valuable for antenatal orientation in Nigeria and contribute to the currently low-resource status of Yorùbá language by improving its availability for NLP tasks.
@@
BibleTTS: BibleTTS [83] is an open speech multilingual dataset containing ten languages of sub-Saharan Africa, viz: Akuapem Twi, Asante Twi, Chichewa, Ewe, Hausa, Kikuyu, Lingala, Luganda, Luo, and Yorùbá. The corpus emanates from Bible recordings provided by the Open-Bible project of Biblica and contains high-quality studio recordings for single speakers, with up to $86$ hours of data per language. Furthermore, TTS models were developed by leveraging the dataset, which ascertains their usability and quality for speech synthesis. The dataset is openly available for commercial and non-commercial use as it is licensed under the CC-BY-SA License, promoting accessibility and further research in text-to-speech for Yorùbá, among other languages captured.
@@
IkiniYorùbá: IkiniYorùbá[27] is a dataset focusing on the cultural aspect of Yorùbá, intending to investigate how well MT models can translate greetings in the language into English. This dataset comprises $960$ parallel sentences of Yorùbá greetings and their English translations. Moreover, it also incorporates the movie transcript subset of the MENYO-20k dataset[13], which contains conversational Yorùbá-English data. The performance of MT models was evaluated on the datasets to enable empirical analysis for related studies in the future. The introduction of the IkiniYorùbá dataset is a valuable contribution to the research community working on Yorùbá NLP.
ÌròyìnSpeech: ÌròyìnSpeech [90] presents high-quality speech corpus exclusively in Yorùbá language, which allows for multipurpose usage involving ASR and TTS tasks. The data collection involves curating 23,000 sentences from news and creative writing domains under an open license (CC-BY-4.0). Also, it included 5,000 sentences on the Mozilla Common Voice platform to crowdsource recordings and validations. The data contributions include 42 hours of in-house recorded speech data by 80 volunteers and an additional 6 hours of validated recordings on Mozilla Common Voice. Ultimately, a high-fidelity single-speaker Yorùbá TTS system was evaluated on the curated speech data, including a baseline word error rate (WER) of $23.8$ for ASR purposes.
@@
MasakhaPOS: MasakhaPOS [54] presents a huge POS dataset available for $20$ diverse African languages, of which Yorùbá language was greatly emphasised. Other languages featured in the dataset include Bambara, Fon, Hausa, Igbo, Luo, Luganda, Akan/Twi, Wolof, isiZulu, isiXhosa, Setswana, Kiswahili, Nigerian Pidgin English, Mossi, Kinyarwanda, Ghomálá, Éwé, chiShona, and Chichewa. Moreover, Yorùbá data was jointly obtained from Voice of Nigeria and Asejere online newspapers, containing a $43,601$ token with an average sentence length of $24.4$. Baseline models using conditional random fields (CRF) and multilingual PLMs such as mBERT and XLM-R were developed to support empirical analysis for related tasks.
MENYO-20k: MENYO-20k [13] presents a multi-domain parallel corpus of Yorùbá-English with clean orthography and standardized train-test splits, with focus on improving the evaluations of MT models on low-resource languages. The dataset’s domain includes texts from news articles, radio and movie transcripts, and TED talks. In addition, special attention was given to the diacritization of Yorùbá texts in the corpus, as it plays an essential role in intelligibility and translation quality in MT NLP tasks. Neural MT models are benchmarked to enable future empirical analysis in similar studies. The dataset is available through a CC BY-NC 4.0 licence.
@@
NollySenti: NollySenti [111] is a parallel multilingual corpus for sentiment classification in five languages spoken in Nigeria, including Yorùbá, Nigerian Pidgin English, Hausa, Igbo, and English. The dataset source reviews from Nollywood movies—movies primarily made in Nigeria—initially in English, with translations into the four Nigerian languages involved. Moreover, the initial collection for the dataset comprises $882$ negative reviews and $1,018$ positive reviews, with an average word length of $20.7$ and $35.0$, respectively, making up two sentiment classification categories. Benchmarks were also developed through traditional machine-learning techniques and pre-trained language modelling to enable empirical analysis for future research in the domain.
Prosodic Read Speech Corpus: This corpus contains a high-quality speech corpus [30] of $7,376$ phrases and sentences in Standard Yorùbá language. A TTS system was developed on the corpus and evaluated using the mean open score and the semantically unpredictable sentence (SUS) test score to support empirical analysis for future related studies. The developed corpus can be used as valuable research material for future studies on Yorùbá speech processing, synthesis, and recognition
TATA: TATA [63] dataset is a large and high-quality multilingual dataset focused on African languages, created from Demographic and Health Surveys (DHS) Program bilingual reports. It consists of $8,700$ examples in nine languages, with Russian as a zero-shot test language. Yorùbá is a core component of these languages; other languages in the dataset include Arabic, English, French, Hausa, Igbo, Portuguese, and Swahili. A transformer-based multilingual PLM was evaluated on the dataset, supporting empirical analysis in future related studies.